1. One integer is 2 more than another. Their product is 35. Find the Integers
2. One integer is 5 less than 4 times another. Their product is 21. Find the integer
3. If 4 is added to the square of an integer the result is 5 less than 10 times that integer. Find the integer

Answer by TimothyLamb(4379)

please post each question separately
below is the answer to the first question:
---
y = x + 2
xy = 35
---
xy = 35
x(x + 2) = 35
xx + 2x - 35 = 0
---
the above quadratic equation is in standard form, with a=1, b=2 and c=-35
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
1 2 -35
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two real roots at:
---
x = 5
x = -7
---
answer:
using either root works for this problem:
x = 5
y = 7
OR
x = -7
y = -5
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php